OpenStudy (anonymous):

Solve the following differential equation: y' - 2xy = 2xe^(x^2)

OpenStudy (anonymous):

Differential equation: \[y' -2xy = 2xe ^{x ^{2}}\] can some one check if this answer is correct? y = x^2*e^\[y = x ^{2}e ^{x ^{2}} + ce ^{x ^{2}}\]

OpenStudy (jamesj):

The integrating factor is u=e^(-x^2) so (uy)' = 2x uy = x^2 + C y = x^2 . e^(x^2) + c.e^(x^2) So, yep. Looks right to me.

OpenStudy (anonymous):

*yeaahhh i finally got a good answer =D

OpenStudy (anonymous):

Thank you somuch JamesJ :)

OpenStudy (jamesj):

np.